Determine the number of carbon and hydrogen types in the following molecules:

1. 3-Pentanol

2. 1-Butanol

3. 2-Pentanol

4. 2,3-Dimethyl-2-butanol

5. 3-Methyl-3-pentanol
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. If you find useful please upvote it. Thanks in advance.

CH3CH2=515H2 Che OM Number of cerbon 25. Type of Hydrogen 32 (1) Primary = 6 (2) Secondary = 5 - 1-butanol 12- on Number of =